twelve divided by the sum of a number and 2 equals the quotient of 4 and the difference of the number and 2. Find the number?
Found 2 solutions by algebrahouse.com, stanbon:
Answer by algebrahouse.com(1079)

"twelve divided by the sum of a number and 2 equals the quotient of 4 and the difference of the number and 2. Find the number?"

x = the number

12/(x + 2) = 4/(x - 2)
12(x - 2) = 4(x + 2) {cross-multiplied, since it was a proportion}
12x - 24 = 4x + 8 {used distributive property}
12x = 4x + 32 {added 24 to both sides}
8x = 32 {subtracted 4x from both sides}
x = 4 {divided both sides by 8}

the number is 4
www.algebrahouse.com

Answer by stanbon(57307)

twelve divided by the sum of a number and 2 equals the quotient of 4 and the difference of the number and 2. Find the number?
--------------
12/(x+2) = 4/(x-2)
----
12(x-2) = 4(x+2)
12x-24 = 4x+8
8x = 32
x = 4
===================
Cheers,
Stan H.